# Review of Machine Learning Approaches for Diagnostics and Prognostics of Industrial Systems Using Industrial Open Source Data

### v2

@@ 3 Detection & Diagnosis
Fault detection and diagnosis are essential components of PHM, tasked with the identification of anomalies and determining failure modes in a timely manner, thereby playing a pivotal role in enhancing system safety and reliability, guaranteeing prompt maintenance interventions, and reducing downtime costs. In this section, we provide a comprehensive overview of the detection and diagnosis challenges encountered in the competitions, as illustrated in Table 3. We sequentially introduce the fundamental competition information, including background, objectives, challenges, and datasets, highlighting the innovative approaches to fault detection and diagnosis problems. A summarization of the competition problems and their datasets is encapsulated in Table 2. Furthermore, we provided a comprehensive summary of the methodologies utilized for detection and diagnosis within the PHM data challenge competition at the end of this section, as outlined in Table 4.
3.1 Challenges
Table 3: Overview of Detection and Diagnosis Challenges in PHM Data Challenge Competitions
Unique Challenges
2021 PHM EU
Missing Data & Noise
Data Imbalance
Limited Data
2022 PHM EU
Missing Data
Data Imbalance
Component-level Prediction
2022 PHM NA Domain Shift
2023 PHM AP
Unknown Fault
Limited Data
2023 IEEE NA
Common Issues: Multiple Fault Modes
Table 3 provides a detailed summary of the detection and diagnosis challenges specific to each competition. While each competition faces unique challenges due to its distinct industry background and objectives, a common issue they all share is the need to accurately classify multiple fault modes. Apart from the common issue, notably, the PHM EU competition in 2021 and the PHM AP in 2023 are both constrained by their small datasets, with the former comprising only 70 training and 29 test samples, and the latter having 177 training and 46 test samples. This paucity of data presents substantial obstacles for training ML models due to limited information from data. Additionally, in the 2021 and 2022 PHM EU competitions, the datasets show a significant imbalance with a higher number of healthy experiments compared to the faulty ones, thus impeding the models’ ability to detect rare fault conditions and challenging their robustness and generalization potential to some degree. In the meantime, these datasets also faced the obstacle of missing data, necessitating imputation techniques for robust data handling. The 2022 PHM NA competition confronts a domain-shift problem, where varying operational conditions and data from different assets result in inconsistent signal distributions, which complicates the task of accurate fault classification. Lastly, the 2023 PHM AP competition includes unexpected fault modes in the testing data, absent from the training set, reflecting the unpredictability of real-world industrial settings and underscoring the necessity for constructing robust and generalized models.
3.2 2021 PHM EU (Manufacturing Production Line)
@@ 3.1 Challenges
In a collaborative effort with the Swiss Centre for Electronics and Microtechnology (CSEM), a data challenge competition was introduced, offering a dataset derived from a real-world industrial manufacturing line dedicated to testing electrical fuses. The objective of this competition is to perform fault identification and classification, root cause analysis, rapid prediction, and system operation parameter identification. The dataset showed eight unique system failure modes, replicated under two distinct operating conditions. However, its scale was rather restricted, comprising merely 70 training and 29 testing samples. The dominant challenge with the training data was its class imbalance—most samples were in healthy condition. This scenario, coupled with the limited volume of data, presented challenges to the data-driven approach.
@@
Against this background, the winning solution was a combination of decision tree algorithms and a propagation system [22]. While the decision trees focused on observational diagnoses, the propagation system skillfully tackled chronological intricacies by infusing a Kalman-style filter. This fusion bolstered the reliability of the outcomes. To address data imbalance, the SMOTE (Synthetic Minority Oversampling Technique) method was utilized. Additionally, [41] incorporated the Leave One Feature Out Importance (LOFO-Importance) package for capturing essential features. Subsequently, they applied linear discriminant analysis (LDA) for dimensionality reduction. During the modeling phase, different ML methods were employed — gradient boosting algorithms such as Extreme Gradient Boosting (XGBoost) and Light Gradient Boosting Machine (LightGBM), LDA classifier, and Gaussian process classifier. Especially for gradient boosting, they used Genetic Algorithms to optimize hyperparameters. A pivotal observation was XGBoost’s superior performance over the other algorithms. Differing from their approach, [3] showcased a rule-based diagnostic technique, comparing its performance with that of decision trees and random forest methods. Additionally, [7] proposed a regularized LSTM for sifting through vital features and then leveraged an ensemble of binary LSTM classifiers for fault detection and classification.
@@
Moreover, the tree-based algorithm XGBoost found favor not just in the aforementioned studies but also in three other distinct research papers [4, 85, 106] outside the competition. The distinctions among these studies were as follows: For instance, [106] relied on a feature importance ranking (FIR) method, targeting enhanced performance and simplification in intricate industrial classification scenarios, whereas [85] focused on the challenge of missing value imputation, harnessing Partial Least Squares (PLS-MV). A significant contribution of Ramezani’s framework was its explainability, achieved by pinpointing key performance indicators for each fault family using the SHAP method [70]. In conclusion, tree-based methods can significantly enhance fault diagnosis, thereby boosting the efficiency and safety of manufacturing production lines.
3.3 2022 PHM EU (Printed Circuit Boards)
@@
The 2022 European PHME Data Challenge, hosted in collaboration with Bitron Spa, spotlights a classification issue within an actual industrial Printed Circuit Board (PCB) production line. The challenge’s objectives span three pivotal sub-tasks: predicting Automatic Optical Inspection (AOI) defect detection based on Solder Paste Inspection (SPI) data, predicting human-made visual inspection labels (OperatorLabel), and finally, predicting the human-assigned repair label (RepairLabel). Task 1 and Task 2 are binary classification problems, whereas Task 3 is a multi-class classification problem in this data challenge. This explains why, in Table 2, the ”Failure Mode” is listed as ”1+1+2”. In Table 2, the ”Operating Condition” is listed as ”1+1+1”, which means each task has its own operating condition. It’s noteworthy that, among the nine datasets discussed in this paper, this dataset stands unique. While others are time series data, this one is parametric data.
@@
Delving into the competition’s outcomes, Gaffet’s team clinched 1st place with an XGBoost method based on encoding and feature engineering [32]. Notably, they further harnessed the SHAP method for model interpretation. Similarly, [102] leveraged two tree-based gradient boosting algorithms, namely LightGBM and XGBoost. Their approach centered on solving classification problems by extracting significant statistical data during the feature engineering phase. The importance of feature engineering was further emphasized by [104]. They introduced a novel statistical feature extraction method coupled with a PinNumber-based technique. This method aimed to compress pin-level data into component-level information. When predicting automatic inspection defects, they integrated a neural network model, factoring in feeding imbalance control to navigate data imbalance challenges. Additionally, a random forest model was developed for both human inspection and repair predictions. Similarly, apart from an extensive focus on feature engineering, [89] applied a multi-layer perceptron (MLP) neural network for defect predictions in automated inspections. For human inspection outcomes, a random forest algorithm was their choice, while decision trees were favored for predicting human repair labels.
@@
Outside of competition, Mirzaei’s team delved deep into the challenge of imbalanced data [75]. Their solution entailed a data-level technique that integrated recursive feature elimination (RFE) for feature selection and oversampling methods. This blend ensured balanced representation for minority classes, enhancing the performance of multiple ML algorithms, from decision trees and random forests to SVMs and 1dCNNs. In a somewhat divergent approach, [58] introduced a new quality management paradigm termed Stream-of-Quality (SoQ) tailored for multi-stage manufacturing processes. By leveraging this dataset, they showcased the effectiveness of their methodology. The implications of their work are vast, offering promising avenues for refining industrial AI algorithms and methodologies in a systematic manner.
3.4 2022 PHM NA (Rock Drill)
@@
Rock drills play a pivotal role in sectors like mining, tunneling, and construction. Due to the potential economic and human costs from work interruptions caused by rock drill faults, ensuring accurate fault diagnosis is imperative. This data challenge tackles the task of fault classification within various rock drill configurations. Challenges arise due to the domain-shift phenomenon, stemming from data sourced from diverse rock drill machines, which results in heterogeneous signal distributions that potentially impair classification accuracy. The dataset [44] encompasses ten distinct fault modes and one health mode, with training, validation, and testing data sizes of 34,045, 3,184, and 16,396, respectively.
@@
In the competition, the champion was clinched by [80], who deployed a hybrid strategy, combining data-driven techniques with various signal-processing methods. Their approach harnessed domain adaptation, metric learning, and pseudo-label-based deep learning to construct an ensemble DL model for comprehensive fault classification. For samples that posed challenges for DL, they deployed signal processing methods like Dynamic Time Warping (DTW), and Cross-correlation, and used SVM for supervised learning. The runner-up solution introduced a data-cropping technique, employing a convolutional neural network (CNN) as a pivotal feature extractor to bridge data length discrepancies. [49] innovatively addressed the domain-shift issue through a domain-adaptation-based scheme that harnessed a domain adversarial learning neural network for extracting domain invariant features while using maximum mean discrepancy (MMD) minimization for bridging distribution discrepancy, and a soft voting ensemble to reduce model uncertainty. Moreover, Minami’s team proposed an ensemble approach, appending specialized models onto a baseline model. This ensemble incorporated domain adaptation strategies to accommodate domain fluctuations. The baseline model employed conventional ML algorithms like SVM, Random Forest, and XGBoost for whole multi-class classification, whereas the specialized sub-models, leveraging CNN for feature extraction and classification, concentrated on binary classifications targeting specific classes that poorly performed on the baseline model [74].
Outside the competition, two notable data-driven strategies emerged.  [65] deployed an end-to-end fault classification framework derived from X-Vectors [95], achieving integrated optimization of both feature extraction and classification phases [65][103] proposed a novel deep forest algorithm that fused multi-grained scanning for feature extraction and a cascade forest for layered predictions to perform failure model classification. In summary, all aforementioned methods emphasize the importance of tackling the domain shift challenge. Utilizing domain adaptation and ML techniques can enhance the accuracy of fault detection and diagnosis in rock drills, ensuring safer and more reliable operations.
@@
3.5 2023 PHM AP (Spacecraft Propulsion System)
@@
The Japan Aerospace Exploration Agency (JAXA) initiated a competition centered on the advancement of PHM technology for spacecraft propulsion systems. The primary objective was to accurately diagnose various states ranging from normal conditions to bubble anomalies, solenoid valve faults, and unknown faults. Simulation data was sourced from four distinct spacecraft, resulting in a dataset comprising 177 training samples and 46 test samples [108]. The paucity of data posed a significant challenge for the data-driven methods.
@@
In the competition, the champions [73], introduced a novel two-step approach. Initially, a similarity measures-based model was proposed for the categorization of data into four distinct states. Subsequently, for data corresponding to the solenoid valve fault, a model incorporating physics-inspired features was employed to pinpoint the fault location and estimate the valve opening ratio. A noteworthy aspect of their methodology entailed the employment of DTW [11] on the training dataset. This step was instrumental in quantifying the variability across various segments of the sensor data. Based on data analysis, authors opted to leverage solely the initial valve opening data segment for feature extraction, thereby mitigating variations arising from individual or sample variations. This strategic refinement substantially improved the data quality, resulting in better model performance in fault detection and classification, with the team achieving 100% classification accuracy in the final competition. Standing in the second position, [61] devised a hybrid approach combining the XGBoost-based method and the rule-based method. While the XGBoost-based approach primarily addressed comprehensive fault classification, the rule-based method was employed to formulate the solenoid valve opening ratio equation. This was accomplished through polynomial fitting, rooted in intrinsic physical characteristics, enabling a precise estimation of the solenoid valve opening ratio. Additionally, [47] utilized the K-NN algorithms to classify faults and pinpoint the location of anomalies. Prior to the classification, a differentiation between normal and anomalous data was executed based on a similarity-based approach. Their estimation metrics for valve opening ratio were hinged upon the similarity of time series waveforms. Moreover, [2] employed an ensemble framework, integrating K-means clustering and decision trees. This model, enriched by domain-specific expertise, exhibited good precision in both anomaly detection and fault diagnosis. Various approaches in the competition underscore the pivotal role of similarity-based methods and the extraction of physical features. These approaches stand out as potent tools in the realm of fault detection and classification, especially when contending with limited datasets.
3.6 2023 IEEE (Gearbox)
Gearbox malfunctions in industrial machinery can result in significant downtime and damage, leading to substantial financial losses. Early detection and accurate diagnosis of these faults are paramount for ensuring the machinery’s safe operation, averting catastrophic failures, and minimizing maintenance costs. The objective of this competition is to develop ML-based models that can efficiently detect faults in the planetary gearboxes of industrial machinery using vibration signals. The dataset [29] spotlights four prevalent sun gear faults: surface wear, chipped teeth, cracks, and missing teeth. Vibration signals have been recorded for a duration of five minutes each, with a sampling rate of 10 kHz, under two distinct operational conditions. Given the expansive data size, totaling 50,000 entries, the potential for deploying DL techniques is evident.
Various data-driven methodologies have been employed in this competition. For instance, [62] utilized ensemble-based CNN techniques to address fault detection via time-series vibration data. Their findings underscored the robust performance of integrating three convolution kernel-based methods such as ROCKET (RandOm Convolutional KErnel Transform) method [23], one-dimensional convolutional neural networks integrated with ResNet, and a fusion of LSTM and Fully Convolutional Network (FCN) in delivering classification results for multivariate time-series data. Similarly, both [91] and [51], designed residual-based CNN models to address fault classification challenges. On the one hand, [91] incorporated the Short-Time Fourier Transform (STFT) [40], transforming frequency domain signals into time-frequency domain signals using Fourier analysis - an invaluable tool for interpreting time-evolving signals. On the other hand, [51] leveraged data augmentation and regularization techniques, enabling model construction with fewer parameters without sacrificing performance quality. In another noteworthy contribution, [96] harnessed a tree classifier to select valuable features from raw vibration signals, subsequently developing a sequential neural network model tailored for the concurrent detection of multiple gear faults.
4 Assessment & Prognosis
This section delves into the health assessment and prognosis problems within PHM, emphasizing their role in evaluating the current health state of industrial machinery and estimating their RUL. Assessment and prognosis are pivotal to implementing predictive maintenance strategies and preemptively addressing system degradation, thereby reducing the incidence of unplanned downtime, extending equipment lifespan, and optimizing maintenance schedules. Similar to the structure of Section 3, we first summarize the challenges in the competition as outlined in Table 5. Subsequently, we detail the specifics of each competition, including the data-driven approaches applied, and provide a consolidated summary of both the problems with their associated datasets and the proposed solutions in Table 6 and Table 7, respectively.
4.1 Challenges
The pivotal challenge across all competitions in health assessment and prognosis is summarized in Table 5. The common challenge focuses on identifying degradation trends and estimating RUL. In the meantime, diverse industrial scenarios present each competition with unique challenges; for instance, the 2018 PHM NA and 2020 PHM EU are both confronted with variability due to differing operational conditions—sensor discrepancies due to varied tool use in the former, and twelve distinct operational conditions influenced by solid ratios and particle sizes in the latter. Similarly, in 2021 PHM NA, the diverse flight envelopes introduce considerable variability into the data. A common thread in the 2018 and 2021 PHM NA competitions is the critical need for models to robustly classify multiple fault modes to enable precise RUL prediction. Complicating this need is the data imbalance evident in the 2018 PHM NA, where rare fault instances are overshadowed by prevalent healthy data, potentially reducing the models’ capabilities for rare fault detection. In the 2019 PHM NA, the limited availability of ultrasonic signal data, particularly for the validation set, poses a risk of model overfitting, potentially reducing the reliability of data-driven prognostics. The 2020 PHM EU is faced with the uncertainty of unknown faults, as the test set includes operational conditions not represented in the training and validation data, a discrepancy that could lead to diminished model performance. These multifaceted challenges necessitate the development of sophisticated, adaptable ML models capable of overcoming data discrepancy, data imbalance, and operational condition variability to deliver accurate and reliable health assessments and prognostic predictions.
Table 5: Overview of Assessment and Prognosis Challenges in PHM Data Challenge Competitions
Unique Challenges
2018 PHM NA
Data Discrepancy & Data Imbalance
Different Operating Conditions
Multiple Fault Modes
2019 PHM NA
Variable Loading Conditions
Limited Physical Information
Limited Data
2020 PHM EU
Different Operation Conditions
Unknown Fault
2021 PHM NA
Multiple Fault Modes
High Variability in Flight Envelopes
Common Issues: Degradation Trend Detection
@@ 4.1 Challenges
4.2 2018 PHM NA (Ion Mill Etching System)

### v3

@@ 3.1 Detection & Diagnosis
In this subsection, we sequentially introduce the fundamental competition information, including background, objectives, challenges, and datasets, highlighting innovative approaches for fault detection and diagnosis problems. A summarization of the competition problems and their datasets is encapsulated in Table 3. Furthermore, we provided a comprehensive summary of the methodologies utilized for detection and diagnosis problems, as outlined in Table 4.
3.1.1 2021 PHM EU (Manufacturing Production Line)
@@
In a collaborative effort with the Swiss Centre for Electronics and Microtechnology (CSEM), 2021 PHM EU offered a dataset derived from a real-world industrial manufacturing line dedicated to testing electrical fuses. The objective of this competition is to perform fault identification and classification, root cause analysis, and system operation parameter identification. The dataset showed eight unique system failure modes under two distinct operating conditions. The primary challenge with the training data was its class imbalance, as the majority of samples represented healthy conditions. Additionally, the dataset was quite small, with only 70 training samples and 29 testing samples. In the fuse test bench dataset, about 10% of the data was missing, and it was not evenly distributed across variables.
@@
Against this background, the winning solution was a combination of decision tree algorithms and a propagation system [23]. While the decision trees focused on diagnosis issue, the propagation system tackled chronology by incoperating a Kalman-style filter. To address data imbalance, the SMOTE (Synthetic Minority Oversampling Technique) method was utilized. Additionally, [45] incorporated the Leave One Feature Out Importance (LOFO-Importance) package for capturing essential features. Subsequently, they applied linear discriminant analysis (LDA) for dimensionality reduction. During the modeling phase, different ML methods were employed — gradient boosting algorithms such as Extreme Gradient Boosting (XGBoost) and Light Gradient Boosting Machine (LightGBM), LDA classifier, and Gaussian process classifier. Especially for gradient boosting algorithms, they used Genetic Algorithms to optimize hyperparameters. A key observation was XGBoost’s superior performance over the other algorithms. Differing from their approach, [3] showcased a rule-based diagnostic technique, comparing its performance with that of decision trees and random forest methods. Additionally, [7] proposed a regularized LSTM for sifting through vital features and then leveraged an ensemble of binary LSTM classifiers for fault detection and classification.
@@
Moreover, the XGBoost algorithm found favor not just in the aforementioned studies but also in three other distinct research papers [4, 91, 112] outside the competition. The distinctions among these studies were as follows: For instance, [112] relied on a feature importance ranking (FIR) method, targeting enhanced performance and simplification in complex industrial classification scenarios, whereas [91] focused on the challenge of missing value imputation, harnessing Partial Least Squares (PLS-MV). A significant contribution of Ramezani’s framework was its explainability, achieved by pinpointing key performance indicators for each fault family using the SHAP method [74].
3.1.2 2022 PHM EU (Printed Circuit Boards)
@@
The 2022 European PHME Data Challenge, hosted in collaboration with Bitron Spa, focused on a classification issue within an actual industrial Printed Circuit Board (PCB) production line. The challenge’s objectives contain three important tasks: (1) Task 1: predicting Automatic Optical Inspection (AOI) defect detection based on Solder Paste Inspection (SPI) data; (2) Task 2: predicting human-made visual inspection labels (OperatorLabel); (3) Task 3 predicting the human-assigned repair label (RepairLabel). Task 1 and Task 2 are binary classification problems, whereas Task 3 is a multi-class classification problem. This explains why, in Table 3, the ”Failure Mode” is listed as ”1+1+2”. In Table 3, the ”Operating Condition” is listed as ”1+1+1”, which means each task has its own operating condition. Meanwhile, the SPI dataset(2022 PHM EU) had missing information in specific fields and 95% of the SPI data was classified as healthy, highlighting a significant imbalance issue.
@@
In 2022 PHM EU, Gaffet’s team got 1st place using XGBoost method based on encoding and feature engineering [32]. Notably, they harnessed the SHAP method for model interpretation. Similarly, [108] leveraged two tree-based algorithms (LightGBM and XGBoost). Their approach centered on solving classification problems by extracting significant statistical data during the feature engineering phase. The importance of feature engineering was further emphasized by [110]. They introduced a novel statistical feature extraction method coupled with a PinNumber-based technique. This method aimed to compress pin-level data into component-level information. When predicting automatic inspection defects, they integrated a neural network model, factoring in feeding imbalance control to navigate data imbalance challenges. Additionally, a random forest model was developed for both human inspection and repair predictions. Apart from feature engineering, [95] applied a multi-layer perceptron (MLP) neural network for defect predictions in automated inspections. For human inspection outcomes, a random forest algorithm was their choice, while decision trees were favored for predicting human repair labels.
@@
Outside of competition, Mirzaei’s team delved deep into the challenge of imbalanced data. They proposed a data-level technique that integrated recursive feature elimination (RFE) for feature selection and oversampling methods. This ensured balanced representation for minority classes, enhancing the performance of multiple ML algorithms, from decision trees and random forests to SVMs and 1dCNNs [80]. Additionally, [62] introduced a new quality management paradigm termed Stream-of-Quality (SoQ) tailored for multi-stage manufacturing processes. By leveraging this dataset, they showcased the effectiveness of their methodology, offering promising avenues for refining industrial AI algorithms and methodologies in a systematic manner.
3.1.3 2022 PHM NA (Rock Drill)
@@
Rock drills play a pivotal role in sectors like mining, tunneling, and construction. Due to the potential economic and human costs from work interruptions caused by rock drill faults, ensuring accurate fault diagnosis is necessary. 2022 PHM NA aimed to address fault classification problem under various product configurations. The main challenges is domain-shift problem, caused by data being collected from different rock drill machines. The dataset encompassed training data from five rock drill machines, validation data from another machine, and test data from two additional machines, highlighting potential domain shift issues between training, validation, and test data. This leads to heterogeneous signal distributions, which can negatively impact classification accuracy. The dataset [48] encompasses ten distinct fault modes and one health mode, with training, validation, and testing data sizes of 34,045, 3,184, and 16,396, respectively.
@@
In the competition,  [86] won the first place. They deployed a hybrid strategy, combining data-driven techniques with various signal-processing methods. Their approach harnessed domain adaptation, metric learning, and pseudo-label-based deep learning to construct an ensemble DL model for comprehensive fault classification. For samples that posed challenges for DL, they deployed signal processing methods like Dynamic Time Warping (DTW), and Cross-correlation, and used SVM for supervised learning. The runner-up solution introduced a data-cropping technique, employing a convolutional neural network (CNN) as a feature extractor to bridge data length discrepancies. [53] innovatively addressed the domain-shift issue through a domain-adaptation-based scheme that harnessed a domain adversarial learning neural network for extracting domain invariant features while using maximum mean discrepancy (MMD) minimization for bridging distribution discrepancy, and a soft voting ensemble to reduce model uncertainty. Moreover, Minami’s team proposed an ensemble approach, appending specialized models onto a baseline model. This ensemble incorporated domain adaptation strategies to accommodate domain fluctuations. The baseline model employed conventional ML algorithms like SVM, Random Forest, and XGBoost for whole multi-class classification, whereas the specialized sub-models, leveraging CNN for feature extraction and classification, concentrated on binary classifications targeting specific classes that poorly performed on the baseline model [79].
@@
Outside the competition, [69] deployed an end-to-end fault classification framework derived from X-Vectors [101], achieving integrated optimization of both feature extraction and classification phases. [109] proposed a novel deep forest algorithm that fused multi-grained scanning for feature extraction and a cascade forest for layered predictions to perform failure model classification.
3.1.4 2023 PHM AP (Spacecraft Propulsion System)
@@
The Japan Aerospace Exploration Agency (JAXA) initiated a competition centered on the advancement of PHM technology for spacecraft propulsion systems. The primary objective was to accurately diagnose various states ranging from normal conditions to bubble anomalies, solenoid valve faults, and unknown faults. Simulation data was sourced from four distinct spacecraft, resulting in a dataset comprising 177 training samples and 46 test samples [114]. The paucity of data posed a challenge for data-driven approaches.
@@
In the competition, the champions [78], introduced a novel two-step approach. Initially, a similarity-based model was proposed for the categorization of data into four distinct states. Subsequently, for data corresponding to the solenoid valve fault, a model incorporating physic-inspired features was employed to pinpoint the fault location and estimate the valve opening ratio. They also deployed DTW [11] on the training dataset which was instrumental in quantifying the variability across various segments of the sensor data. Standing in the second position, [65] devised a hybrid approach combining the XGBoost-based method and the rule-based method. While the XGBoost-based approach primarily addressed comprehensive fault classification, the rule-based method was employed to formulate the solenoid valve opening ratio equation. This was accomplished through polynomial fitting, rooted in intrinsic physical characteristics, enabling a precise estimation of the solenoid valve opening ratio.
Additionally, [51] utilized the K-NN algorithms to classify faults and pinpoint the location of anomalies. Prior to the classification, a differentiation between normal and anomalous data was executed based on a similarity-based approach. Their estimation metrics for valve opening ratio were hinged upon the similarity of time series waveforms. Moreover, [2] employed an ensemble framework, integrating K-means clustering and decision trees. This model, enriched by domain-specific expertise, exhibited good precision in both anomaly detection and fault diagnosis. Various approaches in 2023 PHM AP underscore the importance of similarity-based methods and the extraction of physical features when the dataset size is small.
@@
3.1.5 2023 IEEE (Gearbox)
The objective of 2023 IEEE is to develop ML-based models that can efficiently detect faults in the planetary gearboxes of industrial machinery using vibration signals. The dataset covers four prevalent sun gear faults: surface wear, chipped teeth, cracks, and missing teeth. Vibration signals have been recorded for a duration of five minutes each, with a sampling rate of 10 kHz, under two distinct operational conditions. Given the substantial dataset, comprising 50,000 samples, the deployment of DL techniques is feasible.
In 2023 IEEE. [66] utilized ensemble-based CNN techniques to address fault detection via time-series vibration data. Their findings underscored robust performance of integrating three convolution kernel-based methods such as ROCKET (RandOm Convolutional KErnel Transform) method [24], one dimensional convolutional neural networks (1dCNN) integrated with ResNet, and a fusion of LSTM and Fully Convolutional Network (FCN) in delivering classification results for multivariate time-series data. Similarly, both [97] and [55], designed residual-based CNN models to address fault classification challenges. On the one hand, [97] incorporated the Short-Time Fourier Transform (STFT) [44], transforming frequency domain signals into time-frequency domain signals using Fourier analysis - an useful tool for interpreting time-evolving signals. On the other hand, [55] leveraged data augmentation and regularization techniques, enabling model construction with fewer parameters without sacrificing performance quality. Additionally, [102] harnessed a tree classifier to select valuable features from raw vibration signals, subsequently developing a sequential neural network model tailored for the concurrent detection of multiple gear faults.
3.2 Assessment & Prognosis
Similar to the structure of subsection 3.1, we detail specifics of each competition, including the problems, challenges, and data-driven approaches applied. We also provide a consolidated summary of both the problems with their associated datasets and the proposed solutions in Table 5 and Table 6, respectively.
3.2.1 2018 PHM NA (Ion Mill Etching System)